the perimeter of a rectangle is 68 cm. the length is 4 cm more than its width. find the dimensions of the rectangle
1
Expert's answer
2012-02-21T09:29:50-0500
Let's denote sides of the rectangle as a-length and b-width.
Then perimeter p=2a+2b=68
Also we know that a=b+4 hence
2a+2b=2(b+4)+2b=68

2b+8+2b=68
4b=60
b=15
and hence a=15+4=19